# Research on Superalignment Should Advance Now with Parallel Optimization of Competence and Conformity

### v1

@@ 4 A Novel Conceptual Framework
Figure 4: Overview of our proposed conceptual framework. The superscripts $+$ and -- indicate aligned and unaligned AI, respectively. Left: (1) Mere competence enhancement might cause risky AI behavior. (2) Directly aligning ASI with capability beyond human score is challenging. (3) A training path by alternating optimization of competence and conformity. Right: Three potential paradigms for evaluating the success of superalignment targeting at $\mathbf{x}$ beyond human understanding.
In Sec. 3, we analyzed that existing paradigms are inadequate for achieving superalignment due to the limitations identified. Moreover, they focus on task competence, i.e., deriving better solutions $\mathbf{y}$ for task $\mathbf{x}$ beyond human capability, while failing to address unpredictable risks (producing safe behavior), as the sensitive scenario $\mathbf{x}$ might be unknown. To handle these problems, we discuss the last subargument in Sec. 2.2 here: superalignment should be achieved via simultaneous optimization of task competence and value conformity, as illustrated in Fig. 4. We first introduce two essential principles for superalignment (Sec. 4.1), then describe our conceptual framework for superalignment (Sec. 4.2) and explore potential methods of evaluating success (Sec. 4.3).
@@ 4.1 Two Essential Principles
As defined in Sec. 2.1, the superalignment problem can be transformed into constructing supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$. Since $\mathcal{S}$ is (partially) synthetic, its quality cannot be guaranteed. While there are several limitations in existing paradigms, two key issues stand out: (1) noise in synthetic signals and (2) mode collapse. To mitigate them, we propose two guiding principles for constructing $\mathcal{S}$.
@@
Principle 1: Determining an appropriate capability and capacity gap. Our ultimate goal is to develop a solution $\mathbf{y}$ for tasks $\mathbf{x}$ that require superhuman capacity $\tau\!\gg\!C(\mathcal{H})$ (Sec. 2.2) and capability $U^{*}$ ($\delta\!=\!|U^{*}-U(\mathcal{H})|$ is too large). Large $\tau$ and $\delta$ cause several problems. For Sandwiching, such a hard $\mathbf{x}$ would be beyond human understanding, and candidate solutions generated by $\mathcal{A}$ with sufficient capacity, $C(\mathcal{A})\!=\!\tau$, be highly deceptive, misleading human judges. Our results (Tab. 3.1, Tab. A.4, Sec. A.2.2, Sec. A.3.1) show that capability mismatches between debaters can bias the judge’s decisions in favor of the stronger sneaky debater. For W2SG, when $|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$ is large, imitation of weak supervision occurs. As reported in (Burns et al., 2023), Fig. 5, Tab. A.4 and Tab. LABEL:tab:noise_sciq, such a large gap makes the student $\mathcal{A}_{i+1}$’s performance lag (that’s why we assume Premise 2 in Sec. 2.2). However, a too small gap is also not optimal. Since the initial capacity $C(\mathcal{A}_{1})$ falls far below $\tau$, making the target task $\mathbf{x}$ non-salient in the student models $\mathcal{A}_{1},\mathcal{A}_{2},\dots$, as mentioned in Sec. 3.3 (in this case, $|U^{*}\!-\!U(\mathcal{A}_{i})|\!>\!|C(\mathcal{A}_{i+1})\!-\!C(\mathcal{A}_{i})|$), failing to elicit necessary capability. Therefore, an appropriate gap is needed.
@@
Principle 2: Increasing the diversity of signals. Synthetic data $\mathcal{S}$ would cause significant mode collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024) and self-preference (Wataoka et al., 2024) and bias (Ziems et al., 2024) (for self-enhancement and W2SG), which not only fails to expand the model’s capability but may also degrade it. Instead of relying on one single data source (human or AI), we propose to combine multiple sources to construct the signals, that is, $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{k}\}\bigcup\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$ where $\mathbf{y}_{i}^{k}\!\sim\!\mathcal{A}_{k}(\mathbf{x}_{i})$ and $\mathbf{y}_{i}^{j}\!\sim\!\mathcal{H}_{j}(\mathbf{x}_{i})$. Diverse signal sources (e.g., heterogeneous models or humans with distinct backgrounds) could help mitigate these issues (Du et al., 2024; Chen et al., 2024). Our preliminary results in Fig. LABEL:fig:rewardbench_evaluation_all, Sec. A.2.4 and Sec. A.3.3 demonstrate that noise of synthetic data negatively affects performance while multiple sources would help. Diversity can be achieved by (1) Human-in-the-Loop Feedback (Hejna III & Sadigh, 2023; Retzlaff et al., 2024), (2) Alternative Backbone Models (Kim et al., 2023), and (3) Role-Playing (Shao et al., 2023; Wang et al., 2023b)
4.2 Our Framework
Following the principles above, we illustrate a conceptual framework for superalignment, as shown in Fig. 4 (left).
@@ 4.2 Our Framework
Define $\mathbf{x}^{*}$ as the hard task (e.g., prove the Riemann Hypothesis), and $\mathcal{A}_{1},\!\dots\!,\mathcal{A}_{K}$ as $K$ models with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, as those used in W2SG (Sec. 3.3). The capability gap is defined as $\delta_{i}\!=\!|C(\mathcal{A}_{i+1})-C(\mathcal{A}_{i})|$. Rather than directly requiring $\mathcal{A}_{i}$ to solve $\mathbf{x}^{*}$, we can find or construct a series of predecessor tasks, $\mathbf{x}_{1},\dots,\mathbf{x}_{K-1}$ [^3] Each task may not be a single instance but rather a class of tasks encompassing multiple instances. For clarity, we describe it using a single $\mathbf{x}$ as a representative. in the same domain (like mathematics) with increasing difficulty, e.g., $\mathbf{x}_{1}$ is a basic arithmetic question, $\mathbf{x}_{2}$ is a calculus question, $\mathbf{x}_{K-1}\!=\!$ “Prove Mordell conjecture”. The difficulty gap between $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$ should match the capacity gap $\delta_{i}$, so that task $\mathbf{x}_{i+1}$ is salient in $\mathcal{A}_{i+1}$. Then we propose to conduct a progressive training process, highlighting two key concepts: Gap Annealing and alternating competence enhancement and value alignment. The complete process is described as below:
@@
Step 1: Supervise $\mathcal{A}_{i}$ by $\mathcal{S}_{i}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}$, where $\mathbf{y}_{i}$ could be either a ground truth or synthetic one produced from the previous iteration. As long as $\mathcal{A}_{i+1}$ has not exceeded the human intelligence bound, ground truth remains available.
Step 2: Without safety constraints, $\mathcal{A}_{i+1}$ with higher capability may exhibit harmful behaviors, either proactively (Bengio et al., 2024b) or passively (Chao et al., 2023), defined as $\mathcal{A}_{i+1}^{-}$. Deploy $\mathcal{A}_{i+1}^{-}$ in real or experimental environments, where its potential risks are identified through observation or red-teaming (Perez et al., 2022), uncovering hazardous scenarios $\hat{\mathbf{x}}$. Derive the corresponding safe behaviors $\hat{\mathbf{y}}$ using $\mathcal{A}_{1},\dots,\mathcal{A}_{i}$ and human participants $\mathcal{H}$.
@@
Step 4: Adaptively determine (increase or decrease) the capacity gap $\delta_{i+1}$ according to the real capacity $C(\mathcal{A}_{i+1})$ and utility $U(\mathcal{A}_{i+1})$ evaluated.
This framework offers two advantages. (1) Dynamic and flexible capacity gap. $\delta$ can be adjusted using an annealing process (Fu et al., 2019; Huang et al., 2020). In the early stages, a smaller $\delta$ helps prevent imitation of weak supervision (when noise is high). As the process advances, the gap can be gradually increased to ensure task saliency, assuming $U(\mathcal{A}_{i})$ is sufficiently high and data noise decreases. (2) Exposure to unknown hazardous scenarios. As AI surpasses human capabilities, its potentially harmful behaviors and triggers become unpredictable. Rather than developing a full ASI with extreme risks (Critch & Krueger, 2020; Goldstein et al., 2023) and aligning it afterward, we can expose risk in a controlled environment, where (a) models are not yet capable of causing catastrophic harm, and (b) researchers have time and ability to mitigate them (Cotra, 2021). Note that diversity enhancement approaches can be adopted in each iteration, fostering exploration and safety.
@@ 4.3 Superalignment Evaluation
The evaluation of superalignment is challenging, as the ground truth $\mathbf{y}$ is unavailable and even unassessable. We discuss three potential paradigms, as shown in Fig. 4 (right).
Surrogate Evaluation:  Estimating AI’s performance in real-world cases by that evaluated in simplified cases (Ruan et al., 2024). By allowing AI to operate within these controlled environments, we may examine AI behavior and identify edge cases. For example, an autonomous navigation AI can be tested in a virtual environment simulating diverse traffic, weather, and pedestrian scenarios (Dosovitskiy et al., 2017; Kiran et al., 2021).
Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales (McKenzie et al., 2023). This method continuously evaluates the model as it scales to detect performance shifts. Perez et al. (2023) propose using AI to generate evaluation datasets, identifying unexpected behaviors. Additionally, we suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern, emerging after initial peaks or valleys in tuning.
Peer Review:  Domain experts can qualitatively assess model outputs and reasoning for alignment. Peer review, shown to enhance academic performance (Double et al., 2020), has also been used to mitigate noisy supervision (Lu & Kong, 2023). Similarly, expert panels can evaluate a paper draft on technical accuracy and clarity. Engaging multiple reviewers ensures a more objective assessment of a model’s performance and alignment with human values.
4.4 Potential Implementation
The conceptual framework described in Sec. 4.2 is promising, but challenges remain in implementing it in practice. In this section, we briefly outline some potential directions. For more detail, please refer to Sec. A.1.
Enhancing Data Diversity  Data diversity is essential to avoid mode collapse (Principle 2). Collaboration among AI models with different specialties improves supervision (Guo et al., 2024), while competition exposes alignment weaknesses and solutions (Khan et al., 2024). However, filtering is essential to prevent undesirable behaviors and avoid noisy signals, as discussed in Sec. 3.3 and Sec. 3.1.
Iterative Training Between Teacher and Student  In iterative teacher-student training when $\mathcal{C}(\mathcal{A}_{i-1})<\mathcal{C}(\mathcal{A}_{i})$, a strong teacher $\mathcal{A}^{t-1}_{i}$ generates a solution $\mathbf{y}^{t-1}_{i}$ to supervise a weaker student $\mathcal{A}^{t-1}_{i-1}$, and the supervised weaker student $\mathcal{A}^{t}_{i-1}$ then provides feedback $\mathbf{y}^{t}_{i-1}$ to improve the teacher $\mathcal{A}^{t}_{i}$ (Yuan et al., 2021). This cyclical process can enhance alignment but may face challenges from limited diverse feedback (Sec. 3.3 and Sec. 3.2).
Search-Based Methods Toward Optimal Learning Direction  Search-based methods optimize alignment strategies by exploring possible alignment strategies (Khanov et al., 2024). Techniques like Monte Carlo tree search (undefinedwiechowski et al., 2022) or gradient-free optimization (Larson et al., 2019) can guide promising directions. However, these methods may require high computational resources.
@@ 5 Alternative Views
Superalignment is Unattainable  Absence of ASI and the lack of informed oversight have led some to argue that superalignment is unattainable (Sec.1). Given the current limitations, humans cannot reliably supervise AI systems when their capabilities exceed human expertise, making it difficult to prevent $\mathcal{A}_{\text{ASI}}$ from deceiving its weaker teacher (Burns et al., 2023). Moreover, while existing alignment methods (Zhou et al., 2023; Rafailov et al., 2023; Gulcehre et al., 2023; Xu et al., 2023; Gou et al., 2024) offer limited success in aligning $\mathcal{A}_{\text{ANI}}$, they face scalability and reliability issues.
Superalignment is an Emergent Property  Building on scaling laws (Kaplan et al., 2020; Wei et al., 2023; Nam et al., 2024), one could argue that superalignment might emerge as a byproduct of intelligence expansion rather than requiring proactive intervention. However, concerns about deceptive alignment (Carranza et al., 2023; Hagendorff, 2024) and inverse scaling (McKenzie et al., 2023) persist. As models develop advanced capability, new challenges arise (Taylor et al., 2016; Carlsmith, 2022; Greenblatt et al., 2024b). Relying solely on emergent properties poses risks, including unpredictability in AI behavior.
Our Counterargument  As discussed in Sec. 2.2, we state our position that superalignment is achievable and research on it should advance immediately. Although current paradigms have limitations, delaying research would pose more significant risks (Federspiel et al., 2023; Hendrycks et al., 2023; Bengio et al., 2024a; Greenblatt et al., 2024a). Instead, we propose a simultaneous optimization approach to achieve superalignment, integrating task competence (Motwani et al., 2024) and value conformity (Amodei et al., 2016; Taylor et al., 2016; Olah, 2016).

### v2

@@ 4 Alternating Optimization Framework
Figure 4: Overview of our proposed conceptual framework. The superscripts ++ and - indicate aligned and unaligned AI, respectively. Left: (1) Mere competence enhancement might cause risky AI behavior. (2) Directly aligning ASI with capability beyond human scores is challenging. (3) A training path by alternating optimization of competence and conformity. Right: Three potential paradigms for evaluating the success of superalignment targeting at $\mathbf{x}$ beyond human understanding.
4.1 Feasibility of Superalignment
Based on definitions in Sec. 2, superalignment is achievable if Premises 1 & 2 hold. We show these can be readily satisfied beyond existing paradigms and operationalized via the framework in Sec. 4.3.
Premise 1: Based on the description in Sec. 3, we believe that superhuman capacity, $C(\mathcal{A}_{\text{ASI}})\!\gg\!C(\mathcal{H})$, can be achieved through model/data scaling. The evidence lies in the fact that scaling laws (Kaplan et al., 2020; Gao et al., 2023) still remain effective (OpenAI et al., 2024). It has been shown that roughly every seven months, the complexity of tasks AI can perform (measured by human time required to complete them, called time horizon), doubles (Kwa et al., 2025). Beyond conventional scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), alternative forms of scaling (Xiao et al., 2024; Wu and Tang, 2024), e.g., inference scaling (Pope et al., 2023; Brown et al., 2024), are being discovered to further boost capacity. Despite concerns about data exhaustion, humans continually generate data, especially when assisted by generative AI (Rafiq et al., 2025; Crawl, 2025). These facts strongly support Premise 1. In fact, once $\mathcal{A}$ masters all human data, which is largely true given the data volume of current AI, $C(\mathcal{A})$ already far exceeds $C(\mathcal{H})$.
Premise 2: When $C(\mathcal{A}_{\text{ASI}})$ is obtained, a moderate gap $\delta_{\mathcal{A},\mathcal{H}}\!=\!|\mathbb{D}_{U}\!-\!\mathbb{D}_{C}|$ should be closed without golden signals. While direct access to golden utility $U(\mathcal{H})$ is infeasible, the capability measure $\mathbb{D}_{U}$ can be approximated via surrogate evaluation (see Sec. 4.4). Thus, minimizing $\delta_{\mathcal{A},\mathcal{H}}$ remains feasible in through a long-term developmental path, if high-quality (semi) synthetic supervision signals can be constructed.This capability expansion mirrors teaching a student multiplication by leveraging addition skills, akin to label propagation (He et al., 2019) and easy-to-hard generalization (Sun et al., 2024), where AI can be trained to solve level-5 math problems using only data on easier (level 1–3) ones (Burns et al., 2023).
@@ 4.2 Two Essential Principles
As justified above, the only prerequisite to manifest superalignment is realizable lies in the high-quality supervision signals $\mathcal{S}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$. The three paradigms in Sec. 3 fail since they cannot guarantee such an $\mathcal{S}$, but only offer supervision with limited quality (due to human dependency in Sandwiching) and high noise (collapse and self-preference in Self-Enhancement, and weak imitation in W2SG). To mitigate them, we propose two guiding principles for constructing $\mathcal{S}$.
@@
Principle 1: Determining an appropriate capability and capacity gap. As discussed above, we need to develop solutions for tasks that require superhuman capacity $\Delta^{(C)}\!=\!\mathbb{D}_{C}[C(\mathcal{A}_{\text{ASI}})\!-\!C(\mathcal{H})]$ (Sec. 3) and capability $\delta_{\mathcal{A},\mathcal{H}}\!=\!|\mathbb{D}_{U}[U(\mathcal{A}_{\text{ASI}}),U(\mathcal{H})]\!-\!\Delta^{(C)}|$. With $U(\mathcal{H}),U(\mathcal{A})$ fixed, a large $\tau$ (corresponding to a small $\delta_{\mathcal{A},\mathcal{H}}$), indicates an overly hard task, causing noisy synthetic signals, e.g., low-quality ones from Sandwiching, as shown in our results (Tab. 3.1, Tab. A.9, Sec. A.7.2, Sec. A.8.1), where the capability mismatches between debaters can mislead the judge’s decisions in favor of the stronger sneaky debater. Besides, when $\Delta^{(C)}$ is large, imitation of weak supervision occurs, which makes the student $\mathcal{A}_{i+1}$’s performance lag (negating Premise 2), as reported in Burns et al. (2023), Fig. 5, Tab. A.9 and Tab. LABEL:tab:noise_sciq. However, a $\Delta^{(C)}$ too small is also suboptimal, as the initial capacity $C(\mathcal{A}_{1})$ falls far below $\Delta^{(C)}$, making the target task non-salient in the student models $\mathcal{A}_{1},\mathcal{A}_{2},\dots$, as mentioned in Sec. 3.3, failing to elicit necessary capabilities. Therefore, an appropriate gap is needed.
@@
Principle 2: Increasing the diversity of signals. While not as a standalone solution to value alignment [^6] We acknowledge the extensive literature on value alignment (Anwar et al., 2024; Sorensen et al., 2024; Hadfield-Menell et al., 2016, 2019; Fisac et al., 2020; Barez and Torr, 2023). In contrast, our work formalizes and investigates alignment for superhuman systems where ground truth supervision is infeasible., diversity of supervision signals plays an important role. As analyzed above, the low-quality $\mathcal{S}$ also originates from model collapse (Shumailov et al., 2024; Gerstgrasser et al., 2024), self-preference (Wataoka et al., 2024) and bias (Ziems et al., 2024), even if we find an appropriate gap $\tau$, which not only fails to expand the model’s capability but may also degrade it. Instead of relying on single data source (human or AI), we propose to combine multiple sources to construct the signals, that is, $\mathcal{S}=\{\mathbf{x}_{i},\mathbf{y}_{i}^{k}\}\bigcup\{\mathbf{x}_{i},\mathbf{y}_{i}^{j}\}$ where $\mathbf{y}_{i}^{k}\!\sim\!\mathcal{A}_{k}(\mathbf{x}_{i})$ and $\mathbf{y}_{i}^{j}\!\sim\!\mathcal{H}_{j}(\mathbf{x}_{i})$. Diverse signal sources (e.g., heterogeneous models or humans with distinct backgrounds) could help mitigate these issues (Du et al., 2024; Chen et al., 2024). Our preliminary results in Fig. LABEL:fig:rewardbench_evaluation_all, Sec. A.7.4 and Sec. A.8.3 demonstrate that multiple sources would reduce noise and avoid bias or collapse.
4.3 The ACCO Framework
To operationalize the two principles above, we propose a conceptual Alternating Competence and Conformity Optimization (ACCO) Framework, as shown in Fig. 4 (left).
@@ 4.3 The ACCO Framework
Define $\mathbf{x}^{*}$ as the hard task (e.g., prove the Riemann Hypothesis), and $\{\mathcal{A}_{k}\}_{i=1}^{K}$ as $K$ models with increasing capacities $C(\mathcal{A}_{1})\!<\!\dots\!<\!C(\mathcal{A}_{K})$, where $C(\mathcal{A}_{K})\!=\!C(\mathcal{A}_{\text{ASI}})$, as those used in W2SG (Sec. 3.3). Rather than directly requiring $\mathcal{A}_{i}$ to solve $\mathbf{x}^{*}$, we can find or construct a series of predecessor tasks, $\{\mathbf{x}_{i}\}_{i=1}^{K-1}$ [^7] Each may not be a single but a class of tasks with multiple instances. For clarity, we use a single $\mathbf{x}$. in the same domain (like mathematics) with increasing difficulty, e.g., $\mathbf{x}_{1}$ is a basic arithmetic question, $\mathbf{x}_{2}$ is a calculus question, $\mathbf{x}_{K\!-\!1}\!=\!$ “Prove Mordell conjecture”. The difficulty gap between $\mathbf{x}_{i}$ and $\mathbf{x}_{i+1}$ should match the capacity gap $\Delta^{(C)}_{i}$, so that $\mathbf{x}_{i+1}$ is salient in $\mathcal{A}_{i+1}$. Then we propose a progressive training process, highlighting two keys: Gap Annealing and alternating competence enhancement and value alignment [^8] Our framework is value-type agnostic, as the utility function $U(\mathbf{x})$ can reflect either universal (e.g., fairness, non-discrimination) or culturally grounded values, depending on the supervision signal. This flexibility supports ongoing research on pluralistic alignment Alkhamissi et al. (2024); Anwar et al. (2024).. The complete process is as below:
@@
Step 1: Supervise $\mathcal{A}_{i}$ by $\mathcal{S}_{i}\!=\!\{\mathbf{x}_{i},\mathbf{y}_{i}\}$, where $\mathbf{y}_{i}$ is either a ground truth or synthetic one produced from the previous iteration. When $\mathcal{A}_{i+1}$ is below human intelligence, ground truth remains available.
Step 2: Without value alignment, $\mathcal{A}_{i+1}$ with higher capability may exhibit harmful behaviors, either proactively (Bengio et al., 2024b) or passively (Chao et al., 2023), defined as $\mathcal{A}_{i+1}^{-}$. Deploy $\mathcal{A}_{i+1}^{-}$ in real or experimental environments, where its potential risks are identified through observation or red-teaming (Perez et al., 2022), uncovering hazardous scenarios $\hat{\mathbf{x}}$. Derive safe behaviors $\hat{\mathbf{y}}$ with $\{\mathcal{A}_{k}\}_{k=1}^{i}$ and human $\mathcal{H}$.
@@
Step 4: Adaptively determine (increase or decrease) the capacity gap $\Delta_{i+1}^{(C)}$ according to the real capacity $C(\mathcal{A}_{i+1})$ and utility $U(\mathcal{A}_{i+1})$ evaluated.
This framework offers two advantages. (1) Dynamic and flexible capacity gap. $\Delta^{(C)}$ can be adjusted using an annealing process (Fu et al., 2019; Huang et al., 2020), and determined by search-based optimization (e.g., Monte Carlo tree search (undefinedwiechowski et al., 2022)) to adaptively explore alignment strategies (Khanov et al., 2024). In the early stages, a smaller $\Delta^{(C)}$ helps prevent imitation of weak supervision (when noise is high). As the process advances, the gap can be gradually increased to ensure task saliency, assuming $U(\mathcal{A}_{i})$ is sufficiently high and data noise decreases. (2) Exposure to unknown hazardous scenarios (to support arg.3). As AI surpasses human capabilities $U(\mathcal{A})\gg U(\mathcal{H})$, its potentially harmful behaviors and triggers become unpredictable. Rather than developing a full ASI with extreme risks (Critch and Krueger, 2020; Goldstein et al., 2023) and aligning it afterward, we can expose risk in a controlled environment, gradually increasing its utility $U(\mathcal{A}_{1})\!<\!\cdots\!<\!U(\mathcal{A}_{i+1})$ where (a) models are not yet capable of causing catastrophic harm, and (b) researchers have the time and ability to mitigate them (Cotra, 2021), providing time to correct misaligned behaviors before the model becomes powerful enough to cause catastrophic scenarios $\hat{\mathbf{x}}$ (e.g., a treacherous turn). In each iteration, diversity enhancement approaches can be adopted in each iteration, fostering exploration and safety. This can be achieved by (1) collaborative and competitive interactions among AI (Khan et al., 2024) to uncover failure cases and solutions, (2) human-in-loop feedback (Hejna III and Sadigh, 2023; Retzlaff et al., 2024), (3) alternative backbone models (Kim et al., 2023), or (4) role-playing (Shao et al., 2023; Wang et al., 2023b). More details are presented in Appendix A.4
@@ 4.4 Superalignment Evaluation
The evaluation of superalignment is challenging, as the ground truth $\mathbf{y}$ is unavailable and unassessable. We discuss three potential paradigms, as shown in Fig. 4 (right), which use approximated utility $\hat{U}(\mathbf{y})$ as a proxy for the golden utility $U(\mathbf{y})$ to obtain an empirical estimate of $\mathbb{D}_{U}$.
(1) Surrogate Evaluation:  Estimating AI’s performance in real-world cases by evaluation in simplified cases (Ruan et al., 2024). By allowing AI to operate within controlled environments, we may examine AI behavior. For example, an autonomous navigation AI can be tested in a virtual environment simulating diverse traffic and pedestrian scenarios Dosovitskiy et al. (2017); Kiran et al. (2021). (2) Extrapolation:  Uncovering emergent behaviors and potential misalignments as AI scales McKenzie et al. (2023). This method continuously evaluates AI as it scales. Perez et al. Perez et al. (2023) propose using AI to generate evaluation datasets and identify unexpected behaviors. We also suggest tracking alignment failures when scaling, as misalignments may follow a U-shaped pattern. (3) Peer Review:  Domain experts can qualitatively assess model outputs for alignment. Peer review, shown to enhance academic performance Double et al. (2020), has also been used to mitigate noisy supervision Lu and Kong (2023). This is similar to experts who evaluate a paper draft on technical accuracy and clarity. See Appendix. A.5 for more detailed and formal descriptions.
@@ 5 Alternative Views
While we argue that research on superalignment should begin now (Sec. 2.2), several alternative views persist. We summarize key points below and address them in Appendix A.1. Some claim that superalignment is unattainable, arguing that without informed oversight, supervising more-capable AI systems is infeasible Burns et al. (2023). Others view superalignment as an emergent property of model scaling Kaplan et al. (2020); Wei et al. (2023), yet this overlooks risks like deceptive behavior Carranza et al. (2023) and inverse scaling McKenzie et al. (2023). A third concern is that superalignment is hard to empirically verify, due to the lack of ground truth Yin et al. (2024). We address this by proposing potential evaluation methods in Sec. 4.4. Some argue there is no fundamental difference from current alignment, viewing superalignment as a continuation of standard alignment Fisac et al. (2020); however, it has distinct challenges under insufficient human oversight, detailed in Appendix A.3. Finally, others suggest superalignment can be achieved through inference scaling Muennighoff et al. (2025), though this cannot expand capacity $C(\mathcal{A})$ (Premise 1 in Sec. 3).